# Non-local Metric-Affine Gravity

### v1

@@ V.1 An Example
This is of course tightly related with the fact that two connections are related through a projective transformation and have their symmetric Ricci tensor equal. Furthermore the connections, in this equivalence class, share also the same autoparallel curves up to a reparameterization of the affine parameter Iosifidis:2018zjj .
@@
from which it is clear that the resulting theory is a Vector-Tensor one, where the vector is of geometric origin. Of course, it goes without saying that in general, these excitations extra degrees of freedom will lead to pathologies. However, it should be noted that the additional terms we consider here when added to the quadratic MAG theory could extend the spectrum of healthy theories just like the inclusions of cubic terms in the action, see Bahamonde:2024efl .
@@ VI Applications to Cosmology
For completeness, we will also report the corresponding cosmological equations for the novel theory (56). Let us recall that, in this case, since only the vector part of torsion is excited, in the FLRW spacetime the full torsion is fully described by a single scalar field and the torsion vector takes the form tsamparlis1979cosmological
@@ Acknowledgements
The authors acknowledge the support of Istituto Nazionale di Fisica Nucleare (INFN), Sezioni di Napoli e di Torino, Iniziative Specifiche QGSKY and MOONLIGHT2. This paper is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology).
@@ VIII.2 Variations
With this fact let us proceed to derive the connection field equations for the Theory (54). The variations for the first two lines in (54) are well known, see e.g. Iosifidis:2021tvx , so we will concentrate on the derivation of the new part. Set
@@
From this and using the relations $S_{\mu\nu\alpha}=N_{\alpha[\mu\nu]}$ and $Q_{\alpha\mu\nu}=2N_{(\mu\nu)\alpha}$ it is then clear that we have a dynamical equation for the distortion which we schematically express as $\Box N=N$ as stated.Let us now illustrate how one can derive the metric field equations for the action (11). The $g$-variation first two lines that contain no box operator is well known, see e.g. Iosifidis:2021bad . The lines containing the box operator are trickier so let us elaborate on them in detail. We shall perform the variation on the first term $\sqrt{-g}S_{\mu\nu\alpha}\Box S^{\mu\nu\alpha}$ and of course the same procedure must then be followed for the rest of the terms. The first thing to do is to perform a partial integration

### v2

@@ V.1 An Example
This is of course tightly related with the fact that two connections are related through a projective transformation and have their symmetric Ricci tensor equal. Furthermore the connections, in this equivalence class, share also the same autoparallel curves up to a reparameterization of the affine parameter Iosifidis et al. (2019).
@@
from which it is clear that the resulting theory is a Vector-Tensor one, where the vector is of geometric origin. Of course, it goes without saying that in general, these excitations extra degrees of freedom will lead to pathologies. However, it should be noted that the additional terms we consider here when added to the quadratic MAG theory could extend the spectrum of healthy theories just like the inclusions of cubic terms in the action, see Bahamonde and Gigante Valcarcel (2025).
@@ VI Applications to Cosmology
For completeness, we will also report the corresponding cosmological equations for the novel theory (69). Let us recall that, in this case, since only the vector part of torsion is excited, in the FLRW spacetime the full torsion is fully described by a single scalar field and the torsion vector takes the form Tsamparlis (1979)
@@ Acknowledgements
We thank Friedrich H. Hehl for useful discussions and comments. The authors acknowledge the support of Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Napoli, Iniziative Specifiche QGSKY and MOONLIGHT2, and the Istituto Nazionale di Alta Matematica (INdAM), Gruppo Nazionale di Fisica Matematica (GNFM). This paper is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology).
@@ VIII.2 Variations
With this fact let us proceed to derive the connection field equations for the Theory (67). The variations for the first two lines in (67) are well known, see e.g. Iosifidis (2022a), so we will concentrate on the derivation of the new part. Set
@@
From this and using the relations $S_{\mu\nu\alpha}=N_{\alpha[\mu\nu]}$ and $Q_{\alpha\mu\nu}=2N_{(\mu\nu)\alpha}$ it is then clear that we have a dynamical equation for the distortion which we schematically express as $\Box N=N$ as stated.Let us now illustrate how one can derive the metric field equations for the action (11). The $g$-variation first two lines that contain no box operator is well known, see e.g. Iosifidis (2022b). The lines containing the box operator are trickier so let us elaborate on them in detail. We shall perform the variation on the first term $\sqrt{-g}S_{\mu\nu\alpha}\Box S^{\mu\nu\alpha}$ and of course the same procedure must then be followed for the rest of the terms. The first thing to do is to perform a partial integration